OpenStudy (anonymous):

sin^(1/2)x*cosx-sin^(5/2)x*cosx=cos^3x*sin^(1/2)x

OpenStudy (anonymous):

verify the equation

OpenStudy (dumbcow):

factor out cos x =cosx(sqrt(sin) - sin^2sqrt(sin)) factor out sqrt(sin x) =sqrt(sin)*cosx*(1-sin^2) Note; trig identity : 1-sin^2 = cos^2 =sqrt(sin)*cos^3x

OpenStudy (anonymous):

what happens to the sqrt of sin^5x

OpenStudy (anonymous):

cause you said factor out a sqrt... which i did but then i got lost with that one part of the equation

OpenStudy (dumbcow):

i converted it to sin^2*sqrt(sin) sqrt(x^5) = sqrt(x^4*x) = x^2*sqrtx

OpenStudy (anonymous):

ok gotcha